# Probing Ranking LLMs: Mechanistic Interpretability in Information Retrieval

### v1

@@ 4. Research Questions
4.1. Known features within RankLlama
@@ 4.1.1. MSLR features
We designed sparse probes for each of the selected MSLR features and conducted experiments on each layer of the Llama-2-7b architecture. All experiments were run with both mean and max activation aggregation over the input sequence. Our findings (Figure 2) are categorized into two groups: (i) Features that exhibit a strong fit in certain layers, (ii) Features that do not correlate to the LLM activations in any layer
Positives: Features that achieve an score greater than 0.85 in any particular layer are considered strong positives. This means it is highly likely that the activation has extracted the particular feature within the neuron. We found that several metrics, namely covered query term number, covered query term ratio, min of term frequency, mean of stream length normalized term frequency, variance if tf*idf and min of tf*idf, frequently exhibit low Mean Square Errors and high scores. This indicates that there exist MLP neurons within the investigated layers that perform feature extractions similar to these MSLR features. Additionally, we observed that the scores obtained with mean and max activation aggregation are quite similar, except minor exceptions, suggesting either aggregation method is usable.
Negatives: Certain features failed to achieve an score greater than 0.1 in the final layer, often achieving a highly negative . This indicates that the LLM does not consider these features important in their current form. Features from the MSLR set that fell into this category include: sum and max of tf*idf; sum and max of stream length normalized term frequency, max and variance of term frequency, stream length and sum of term frequency. Thus, while mean of stream length normalized term frequency is something the LLM highly seeks, it does not seek for the sum and max of normalized term frequency at all!
@@ 4.1.2. Traditional similarity metrics
We also probe the LLM architecture for non-semantic statistical query-document distance metrics like tf*idf cosine score, euclidean score, manhattan score, Kullback-Leibler divergence score and Jensen-Shannon divergence score. Our motive behind this is to identify if Llama includes components that mimic statistical similarity measures. Results for the probe on RankLlama-7b are visualized in Figure 3. We observe that statistical score based methods do not correlate well with neuron activations. Out of the feature examined, euclidean score shows the highest correlation with reaching , dipping before the final layer. Given its well known success in the past, it is ironic that the tfi*df cosine score between the query and document shows the least correlation when probed for, within RankLlama activations.
@@ 4.1.3. BERT and T5 scores
BERT and T5 are neural models widely used for retrieval and reranking tasks before the advent of LLMs. LLMs are much larger in size compare to BERT models. As a result it is of interest to see if BERT and T5 subnetworks are present within RankLlama activations. We design probes to mine for the cosine distance between the BERT and T5 embeddings of the query and the document. Our observations (presented graphically in supplementary material) show that both BERT and T5 obtain moderate scores in our probes, reaching and on RankLlama 7b and 13b respectively. The BERT and T5 scores follow each other across the LLM layers. This indicates that the LLM potentially does not encode BERT and T5 subnetworks as-is. We compare our findings to related works on BERT probing in the Discussion (Section 6).
@@ 4.2. Feature Groups within RankLlama
In the previous subsection, we found that several MSLR features like covered query term number, covered query term ratio, mean of stream length normalized term frequency, min of tf/tf*idf and variance of tf*idf appear to be modelled within LLM activations, especially within the later layers. These features also seem to be a fair choice to model relevance based on intuition. However, these features might not be present directly but in combination with one another or individually with different exponents. To test for this possibility, we probe for combinations of these features within different layers of the LLM and find various combinations of covered query term ratio, mean of stream length normalized term frequency and variance of tf*idf. We find strong indication for their presence within the LLM activations of the later layers of the LLM. For example, if QTR represents covered query term ratio, STF represents mean of stream length normalized term frequency and VTFIDF represents variance of tf*idf, we find high scores on average for all of QTR+STF, QTR+VTFIDF, STF+VTFIDF, QTR+STF+VTFIDF, QTR*STF, STF*VTFIDF, QTR*VTFIDF, and QTR*STF*VTFIDF when probing the last layer of the LLM. This potentially indicates that the LLM has over the course of layers learned some representation of (QTR+STF+VTFIDF)k. In Figure 4, we show the probing performance of a sum of those three features and their exponents, relative to BM25. We observe that the sum of this feature group, its square, as well as its cube consistently show a strong correlation to neuron activations within RankLlama.
4.3. Out-of-distribution queries and documents
We know that RankLlama 7b and 13b have been obtained by fine-tuning Llama-2 7b and 13b using LoRa on the MS MARCO dev set (Ma et al., 2023). It is of interest to see if the fine-tuned LLM extracts different features from in-distribution and out-of-distribution query-document pairs. As a result, we try probing RankLlama with query-document pairs from 2 others datasets. Firstly, we use the BEIR Scidocs dataset (Thakur et al., 2021), a dataset comprising of scientific documents and queries. We select 200 queries and use bm25 to retrieve the top 1000 documents for our set of queries. We then probe the MS MARCO fine-tuned RankLlama with BEIR activations. We repeat the process with the SoDup dataset (Lambert et al., 2023), which contains a question from stack overflow as query and a list of other relevant questions from Stack overflow as potential duplicate candidates. For each question, we select known duplicate questions from the dataset and treat them as relevant documents. We again pick 200 queries from this dataset for our out-of-distribution probing experiments. We probe for each of the 19 MSLR features on both the RankLlama model sizes and compare probing results between in-distribution and the out-of-distribution datasets. We report our observation in Figure 5. For a given feature, e.g: mean of tf, we see that our probes can not find any activations representing the feature in RankLlama 7b, both with in-distribution and out-of-distribution datasets. However, when probing RankLlama 13b, while out-of-distribution activations do not capture this feature, it gets a hit within in-distribution activations.
We see some variance in results between RankLlama 7b and 13b in this set of experiments. The probes seem to fare similarly on most features between in-distribution and out-of-distribution data on rankLlama 7b. The only exception being sum of stream length normalized tf, which has a negative score for both in and out distribution data, making it’s magnitude insignificant. However, the probes seem to fetch different results between in and out-of-distribution data for RankLlama 13b. It in particular finds probes strongly correlated to stream length, sum and mean of tf, and mean of tf*idf in the in-distribution data, but not for the out-of-distribution data. A likely reason for this might be that RankLlama 13b has overfit to the MS MARCO dev set and hence seeking features like stream length which are known to not generalize for mapping query-document relevance (Belinkov, 2022).
5. Validating Probes
While probing primarily serves to generate hypotheses, these can be independently validated using various methods. In this section, we aim to verify the findings from our probing experiments, particularly focusing on neurons that showed strong correlations with probed features within the activations of the RankLLaMA model. We concentrate on the probing patterns in the last layer, as this layer is directly responsible for ranking decisions. To ensure the robustness of these findings, we employ feature attribution analysis using Shapley values (Sundararajan and Najmi, 2020). Shapley values allow us to quantitatively assess the contributions of different neurons and neuron groups to the model’s predictions, thus providing an indirect validation of the hypotheses generated through probing. The approach internally conducts an ablation study of measuring each neuron’s individual contribution to the ranking decision and averages them based on some game theoretic principles(Chowdhury et al., 2024).
We experiment on both RankLLaMA models and focus on the features found to have a strong correlation () within neurons of the final layer. We work with the same MS MARCO query-documents pairs as used in the probing experiments and use the change in NDCG score on addition of a feature as its value function for feature attribution computation (Lundberg and Lee, 2017).
We run our probing validation experiments for a hundred query-document pairs on each RankLlama model and compute Shapley values of each neuron in the last layer of the LLMs. We report average attributions across neuron groups containing 1 neuron, 2-3 neurons and 4+ neurons. Our results can be found in Table 1. We find that the average feature attribution of neurons in our identified neuron groups is within the 99th percentile in 79 out of the 100 inputs in RankLlama 7b and 84 out of the 100 inputs in RankLlama 13b. This indicates that the identified neurons within the LLM are instrumental in making the ranking decision, thus validating the location of our probes.
| Model | 1-neuron | 2-3 neurons | 4+ neurons |
| RankLlama 7b | 84 | 79 | 74 |
| RankLlama 13b | 87 | 82 | 78 |
@@ 6. Discussion
Comparing Notes with Previous Probing Work: A number of studies have probed BERT and T5 with the aim of understanding if they encode concepts like term frequency and inverse document frequency. Formal et al. (Formal et al., 2021) study the matching process of ColBERT (Khattab and Zaharia, 2020) and conclude that the model is able to capture a notion of term importance and relies on exact matches for important terms. This is in agreement with our findings, where we find strong correlations to term matching features like covered query term number/ratio and stream length normalized term frequency within LLM activations. In a later work, Formal et al. (Formal et al., 2022) conduct a study to measure the out-of-domain zero-shot capabilities of BERT/T5 models in lexical matching on the BEIR dataset, and show that these models fail to generalize lexical matching in out of domain datasets or terms not seen at training time. This finding continues for RankLlama as well, where the model is unable to generalize lexical matching to terms not seen beforehand (Figure 5, RankLlama 13b OOD vs In-distribution probing discrepencies).
@@
Limitations of Probing: While probing techniques are widely used to understand the internal workings of LLMs, there are several limitations of using probing for this purpose. (1) Probing techniques reveal correlations between specific features and neuron activations, but not a causal analysis of the decision making process. (2) The insights gained from probing are heavily dependent on the dataset used for probing. If the dataset is not representative of the model’s typical input, the results may not accurately reflect the model’s general behavior. (3) LLMs often rely on complex interactions between multiple features. Probing techniques may fail to capture these interactions, leading to an incomplete understanding of how the model makes decisions. (4) Probing can identify features but may mislead regarding their role or significance in the model’s decision-making process. However, many of the limitations of probing can be addressed by employing various techniques, such as creating balanced probing datasets and independently validating probing results through methods like ablation studies and feature attribution analyses. Ultimately, probing is an effective tool for generating hypotheses about the internal workings of LLMs, provided that these hypotheses are subsequently validated through independent methods.
@@ 7. Conclusion and Future Work
In this study, we probed the Llama-2 7b and 13b large language models, each fine-tuned for passage reranking, to look for a subset of features from the MSLR dataset. These human-engineered features, which include query-only, document-only, and query-document interaction features, are recognized as significant in ranking tasks. Using a layer-wise probe, we discovered that the activations in most layers could accurately replicate select features such as min tf*idf, covered query term number/ratio, min of term frequency, and mean of stream length normalized term frequency. This suggests that the fine-tuned Llama network deems these features important. Conversely, we found no correspondence for features like sum and max of tf*idf and sum and max of term frequency in the probed neurons. This indicates the absence of these features in RankLlama’s neural feature extractors. We also reported observations on feature groups that show a strong correlation with RankLlama neurons and hypothesized some abstract features mined by the LLM. Finally, we compared correlated features between in-distribution and out-of-distribution datasets and encountered a case of the LLM’s apparently overfitting during fine-tuning. These findings enhance our understanding of open-source ranking LLMs.
We used ridge regression based linear probes in this study. While linear probes are simple and interpretable, it is unable to disentangle features distributed features that combine in a non-linear way. In the future, it would be interesting to use non-linear probes, such as decision trees or simple neural nets to disambiguate non-linear concepts. The long-term objectives of this endeavor include: (i) identifying potential modifications to existing MSLR features, such that they align further with LLM activations, (ii) deciphering and reverse-engineering segments of the LLM that do not correspond to recognized MSLR features, and (iii) ultimately cataloging all features deemed significant by LLMs and plugging them back into simpler statistical models to enhance the performance and interpretability of statistical ranking models.
These directions will help further explain the underlying mechanisms of ranking LLMs and improve their interpretability.

### v2

@@ 4. Research Questions
4.1. Statistical features within RankLlama2-7b
@@ 4.1.1. MSLR features
For the first set of experiments, we designed sparse probes for each of the selected MSLR features and conducted experiments on each layer of the RankLlama2-7b architecture. All experiments were run with both mean and max activation aggregation over the input sequence. Our findings (Figure 2) are categorized into two groups: (i) Features that exhibit a strong fit in certain layers, (ii) Features that do not correlate to the LLM activations in any layer
Positives: Features that achieve an score greater than 0.85 in any particular layer are considered strong positives. This means it is highly likely that the activation has extracted the particular feature within the neuron. We found that seven metrics, namely Min of TF, Min(TF/L), Min of TF*IDF, covered QT Number, covered QT Ratio, Mean(TF/L), and variance of TF*IDF frequently exhibit low Mean Square Errors and high scores. This indicates that there exist MLP neurons within the investigated layers that perform feature extractions similar to these MSLR features. Additionally, we observed that the scores obtained with mean and max activation aggregation are comparable, except minor exceptions, suggesting either aggregation method is suitable.
Negatives: Certain features failed to achieve an score greater than 0.1 in the final layer, often achieving a highly negative . This indicates that the LLM does not consider these features important in their current form. The 10 features from the MSLR set that fell into this category include: Sum and Max of TF*IDF, Sum(TF/L) and Max(TF/L), Max and variance of TF, stream length and Sum of TF. Thus, while Mean(TF/L) is something the LLM highly seeks, it does not seek for the sum and max of normalized term frequency at all!
@@ 4.1.2. Traditional similarity metrics
We also probe the LLM architecture for non-semantic statistical query-document distance metrics like tf*idf cosine score, euclidean score, manhattan score, Kullback-Leibler divergence score and Jensen-Shannon divergence score. Our motive behind this is to identify if the LLMs include components that mimic statistical similarity measures. Results for the probe on RankLlama2-7b are visualized in Figure 3. We observe that statistical score based methods do not correlate well with neuron activations. Out of the feature examined, euclidean score shows the highest correlation with reaching , dipping before the final layer. Given its well known success in the past, it is ironic that the tfi*df cosine score between the query and document shows the least correlation when probed for, within RankLlama activations.
@@ 4.1.3. BERT and T5 scores
BERT and T5 are neural models widely used for retrieval and reranking tasks before the advent of LLMs. LLMs are much larger in size compared to BERT models. As a result it is of interest to see if BERT and T5 subnetworks are present within the studied LLM’s activations. We design probes to mine for the cosine distance between the BERT and T5 embeddings of the query and the document. Our observations show that both BERT and T5 obtain moderate scores in our probes, reaching and on RankLlama2-7b and 13b respectively. The BERT and T5 scores follow each other across the LLM layers. This indicates that the LLM likely does not encode BERT and T5 subnetworks as-is. We compare our findings to related works on BERT probing in Section 5.
4.2. Comparing Ranking LLMs
We next explore whether different ranking LLMs encode similar or distinct Information Retrieval features. To investigate this, we extend our analysis beyond the RankLlama2-7b model and compare it to the two additional fine-tuned LLMs, RankLlama3-8b and RankPythia-6.9b, as well as the RankLlama2-13b model. We fine-tune these models under an identical passage reranking framework, with an identical dataset and loss function, to ensure a fair and consistent comparison.
Once the models are fine-tuned, we conduct the same probing experiments to analyze how well each ranking LLM encodes the known human-engineered and semantically meaningful attributes relevant to ranking tasks. The results of these probing experiments are summarized and visualized in a spider chart (Fig. 4), allowing for an intuitive comparison of feature encoding across the four models. It highlights the similarities and differences in the latent features captured by LLMs with varying architectures and sizes, shedding light on the generalizability and architecture-specific tendencies of ranking-focused fine-tuning. The figure depicts the score for an IR feature, when probed in a particular LLM, and lies between .
All four ranking LLMs demonstrate broadly comparable patterns for most features, particularly for the Min metrics (e.g., Min of TF, Min of TF*IDF), where they uniformly show top scores (all 1.0). They also appear closely aligned on coverage-based features—such as Covered Query Term Number and Ratio—with each scoring in the high 0.95–0.98 range. By contrast, the greatest differences emerge in TF-based statistics like Max of Term Frequency and Variance of TF/TF*IDF, where RankPythia tends to exhibit slightly higher peaks (e.g., a larger for Max of TF) than the RankLLaMA models. Additionally, all LLMs show near zero performance in encoding Sum of Stream Length Normalized TF, suggesting similarity in how these models handle term-frequency weighting. Overall, the four models encode IR features in a largely similar manner, indicating that our probing findings are generalizable and not LLM specific.
4.3. Out-of-distribution queries and documents
All four of the LLMs we study were obtained by fine-tuning the base models using LoRa on the MS MARCO dev set (Ma et al., 2023). It is of interest to see if during inference the fine-tuned LLM extracts different features from in-distribution and out-of-distribution query-document pairs. We try probing all the studied LLMs with query-document pairs from two others datasets. First, we use the BEIR Scidocs dataset (Thakur et al., 2021), comprising scientific documents and queries. We select 200 queries and use bm25 to retrieve the top 100 documents for each. We then probe the MS MARCO fine-tuned LLMs with BEIR activations. We repeat the process with the SoDup dataset (Lambert et al., 2023), which contains a question from StackOverflow as query and a list of other relevant questions from StackOverflow as potential duplicate candidates. For each question, we select known duplicate questions from the dataset and treat them as relevant documents. We again pick 200 queries from this dataset for our out-of-distribution probing experiments. We probe for each of the 19 MSLR features on all the fine-tuned ranking LLMs and compare probing results between in-distribution and the out-of-distribution datasets. We show our comparision of the probing results of RankLlama2-7b vs RankLlama2-13b in Figure 5, reporting the score of each feature in the last layer of the respective LLM. For the feature mean of tf we see that our probes cannot find any activations representing the feature in RankLlama 7b, both with in-distribution and out-of-distribution datasets. However, when probing RankLlama 13b, while out-of-distribution activations do not capture this feature, it gets a hit within in-distribution activations.
We see some variance in results between RankLlama2-7b and RankLlama2-13b in this set of experiments. The probes seem to fare similarly on most features between in-distribution and out-of-distribution data on RankLlama2-7b. The only exception being sum of stream length normalized tf, which has a negative score for both in and out distribution data, making its magnitude insignificant. However, the probes seem to fetch different results between in and out-of-distribution data for RankLlama2-13b. It in particular finds probes strongly correlated to stream length, sum and mean of tf, and mean of tf*idf in the in-distribution data, but not for the out-of-distribution data. A likely reason for this might be that RankLlama 13b has overfit to the MS MARCO dev set and is hence seeking features like stream length which are known to not generalize for reflecting query-document relevance (Belinkov, 2022).
We compared the probing performance of in-distribution and out-of-distribution query-document pairs in RankLlama3-8b and RankPythia-6.9b, observing similar performance for both types of inputs, consistent with the results observed for RankLlama2-7b in Figure 5a. The graphs depicting these probing results are not included due to space constraints.
@@ 4.4. Feature Groups within RankLlama
In the previous subsection, we found that several MSLR features (covered QT Number, covered QT Ratio, mean(TF/L), min TF, min of TF*IDF, Min(TF/L) and variance of TF*IDF) appear to be modeled within LLM activations, especially within the later layers. These features also seem to be a fair choice to model relevance based on intuition. However, these features might not be present individually but in combination with one another or individually with different exponents. To test for this possibility, we probe for combinations of these features within different layers of the LLM and find various combinations of covered query term ratio, mean of stream length normalized term frequency and variance of tf*idf. We find strong indication for their combined presence within the LLM activations of the later layers of the LLM. For example, if QTR represents covered query term ratio, STF represents mean of stream length normalized term frequency and VTFIDF represents variance of tf*idf, we find high scores on average for all of QTR+STF, QTR+VTFIDF, STF+VTFIDF, QTR+STF+VTFIDF, QTR*STF, STF*VTFIDF, QTR*VTFIDF, and QTR*STF*VTFIDF when probing the last layer of the LLM. This potentially indicates that the LLM has over the course of layers learned some representation of (QTR+STF+VTFIDF)k. In Figure 6, we show the probing performance of a sum of those three features and their exponents, relative to BM25 in RankLlama2-7b and 13b models. We observe that the sum of this feature group, its square, as well as its cube consistently show a strong correlation to neuron activations within RankLlama. This analysis provides a foundation for future work to uncover and investigate complex groups of features encoded within LLM activations. By highlighting how various ranking architectures capture and represent different IR features, it opens avenues for deeper exploration into latent feature interactions and the mechanistic underpinnings of ranking-focused language models.
@@ 5. Discussion
Comparing Notes with Previous Probing Work: A number of studies have probed BERT and T5 with the aim of understanding if they encode concepts like term frequency and inverse document frequency. Formal et al. (Formal et al., 2021) study the matching process of ColBERT (Khattab and Zaharia, 2020) and conclude that the model is able to capture a notion of term importance and relies on exact matches for important terms. This is in agreement with our findings, where we find strong correlations to term matching features like covered query term number/ratio and stream length normalized term frequency within LLM activations. In a later work, Formal et al. (Formal et al., 2022) conduct a study to measure the out-of-domain zero-shot capabilities of BERT/T5 models in lexical matching on the BEIR dataset, and show that these models fail to generalize lexical matching in out-of-domain datasets or terms not seen at training time. This finding continues for RankLlama as well, where the model is unable to generalize lexical matching to terms not seen beforehand (Figure 5, RankLlama 13b OOD vs In-distribution probing discrepencies).
@@
Validating Probing Results While probing techniques are widely used to understand the internal workings of LLMs, there are several limitations of using probing for this purpose. (1) Probing techniques reveal correlations between specific features and neuron activations, but not a causal analysis of the decision making process. (2) The insights gained from probing are heavily dependent on the dataset used for probing. If the dataset is not representative of the model’s typical input, the results may not accurately reflect the model’s general behavior. (3) LLMs often rely on complex interactions between multiple features. Probing techniques may fail to capture these interactions, leading to an incomplete understanding of how the model makes decisions. (4) Probing can identify features but may mislead regarding their role or significance in the model’s decision-making process. However, many of the limitations of probing can be addressed by employing various techniques, such as creating balanced probing datasets and independently validating probing results through methods like ablation studies and feature attribution analyses.
To validate our probing results, we analyzed neurons in the final layer of RankLLaMA models that exhibited strong correlations () with probed features, as this layer directly influences ranking decisions. Using Shapley values (Datta et al., 2016; Lundberg and Lee, 2017; Chowdhury et al., 2024) for feature attributions we assessed the contributions of individual neurons and neuron groups to the ranking predictions, leveraging a value function based on changes in NDCG scores for MS MARCO query-document pairs. Probing validation was conducted on 100 query-document pairs for each RankLLaMA model, and we computed average attributions across neuron groups of varying sizes. Results showed that identified neuron groups ranked within the 95th percentile in 79 out of 100 cases for RankLLaMA 7B and 84 out of 100 for RankLLaMA 13B, confirming that these neurons are instrumental in ranking decisions and validating the location of the probes.
@@ 6. Conclusion and Future Work
In this study, we probed four different large language models, each fine-tuned for passage reranking, to look for a subset of features from the MSLR dataset. These human-engineered features, which include query-only, document-only, and query-document interaction features, are recognized as significant in ranking tasks. Using a layer-wise probe, we discovered that the activations in most layers could accurately replicate select features such as Min TF*IDF, covered QT Number/Ratio, Min TF, Min(TF/L), and Mean(TF/L). This suggests that the fine-tuned Llama network deems these features important. Conversely, we found no correspondence for features like sum and max of tf*idf and sum and max of tf in the probed neurons. This indicates the absence of these features in RankLlama’s neural feature extractors. We also reported observations on feature groups that show a strong correlation with RankLlama neurons and hypothesized that some abstract features are mined by the LLM. We also observed that different LLM architectures, when fine-tuned on the same dataset with the same loss function, exhibit similar latent features when probed. Finally, we compared later features encoded between in-distribution and out-of-distribution datasets and encountered a case of the LLM’s apparently overfitting during fine-tuning. These findings enhanced our understanding of ranking LLMs and gave us generalizable insights for the design of more effective and transparent ranking models. The long-term objectives of this endeavor include: (i) identifying potential modifications to existing MSLR features, such that they further alight with LLM activations, (ii) deciphering and reverse-engineering segments of the LLM that do not correspond to recognized MSLR features, and (iii) ultimately cataloging all features deemed significant by LLMs and plugging them back into simpler statistical models to enhance the performance and interpretability of statistical ranking models. These directions will help further explain the underlying mechanisms of ranking LLMs and improve human trust into them.
Acknowledgments
This work was supported in part by the Center for Intelligent Information Retrieval. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.